A Mercedes of mass 200 kg is traveling west at 20 m/s and collides with a 400 kg smart car that is traveling east at 30 m/s. If after the collision, the smart car rebounds west at 15 m/s, what is the velocity and direction of the Mercedes AFTER the collision?


1
Expert's answer
2021-02-18T18:46:12-0500
"mv_1-Mu_1=-mv_2+Mu_2\\(200)(20)-(400)(30)=-(200)v_2+(400)15\\v_2=70\frac{m}{s}"

Westwards.